OpenStudy (sillybutt100):

How do I solve this What is the product of (6)sqrt7 (2) sqrt21? Simplify if possible I want to know how to do this. The equation is also in the comments

OpenStudy (jhannybean):

is it (6 sqrt(7))^2? or 6 (sqrt(7))^2...

OpenStudy (sillybutt100):

\[6\sqrt{7} \times 2 \sqrt{21}\]

OpenStudy (sillybutt100):

Sorry I fixed it.

OpenStudy (bahrom7893):

18*sqrt(7*7*7*7)

OpenStudy (bahrom7893):

does that help?

OpenStudy (jhannybean):

you can combine both of these. multiplying 2 square roots \(\large \sqrt{a}*\sqrt{b} = \sqrt{ab}\)

OpenStudy (mertsj):

\[6\sqrt{7}\times 2\sqrt{21}=6\times2 \times \sqrt{7 \times 3 \times 7}\]

OpenStudy (jhannybean):

Does whatever these helpers mentioned help you understand how it works?...

OpenStudy (sillybutt100):

sort of hold on

OpenStudy (sillybutt100):

So far I have \[12*\sqrt{147}\]

OpenStudy (jhannybean):

and 147 = 21 * 3 = 7 * 3 * 3\[\large \sqrt{147} = \sqrt{7*\color{blue}{3*3}} \] see how we can pull out the 3's?

OpenStudy (sillybutt100):

not really

OpenStudy (jhannybean):

for every pair of 2 numbers (of the similar kind) inside a square root, you can pull one out. Therefore if we have 2 3's inside the square root, we can pull one out. This will leave us with \[\large 3\sqrt{7}\]

OpenStudy (sillybutt100):

That is it? It is solved?

OpenStudy (jhannybean):

I was just explaining how you simplify stuff inside the square root.

OpenStudy (jhannybean):

Now you can refer back to what @Mertsj has shown up there, and use this concept of pulling out pairs of numbers inside the square root to solve your problem

OpenStudy (sillybutt100):

Thanks

OpenStudy (sillybutt100):

is the answer \[7\sqrt{3}\]

OpenStudy (mertsj):

no

OpenStudy (mertsj):

\[6 \times 2 \sqrt{7 \times7 \times 3}=12\times 7 \sqrt{3}=84\sqrt{3}\]

OpenStudy (sillybutt100):

Oh I left out the 12 when multiplying it with the 7

OpenStudy (sillybutt100):

@Jhannybean When it says "What is the exact value of the expression" then has an equation like the last one what does it want me to do?